Charlotte (age 40) is a surviving spouse and provides all of the support of her four minor children who live with her (all are under age 16). She also maintains the household in which her parents live and furnished 60% of their support. Besides interest on City of Miami bonds in the amount of $5,500, Charlotte's father received $2,400 from a part-time job. Charlotte has a salary of $80,000, a short-term capital loss of $2,000, a cash prize of $4,000 from a church raffle, and itemized deductions of $10,500. Click here to access the standard deduction table to use, if required. a. Compute Charlotte's taxable income. $ b. Using the Tax Rate Schedules (click here), tax liability for Charlotte is $ for 2019. c. Compute Charlotte's child and dependent tax credit.
Answer #1

Ans:

A). Charlotte Taxable income=

Salary= 80,000

Short term capital loss= -2,000

Cash prizes= 4,000

AGI= 82,000

Less: Gretaer of standard or itemzied deduction= 24,400

Taxable income= 57,600

B). Tax liability (surviving spouse)= (1940)+12% (57,600-19,400)

=1,940+12% (38,200)

= 1,940+4,584

= 6,488

C). Child and dependent tax credit= 2000*4+500*2

= 9,000